the area of a square is 64 square inches. what is the length of each side.
Answer by nerdybill(7384)

the area of a square is 64 square inches. what is the length of each side.
.
Let x = length of one side of the square
then
x*x = 64
x^2 = 64
x = sqrt(64)
x = 8 inches